### v2

@@ B.2 Proof of Lemma 13
We use ${\mathcal{A}}$ to denote the greedy subroutine in the Random Subtree algorithm. The tie-breaking rule of ${\mathcal{A}}$ is local by definition, and it remains to show that it is also non-increasing. For each node $x$ in the HST, denote $A_{\ell}(x)$ as the $\ell$-th ancestor of $x$ and $T_{\ell}(x)$ as the subtree rooted at $A_{\ell}(x)$. That is, $A_{0}(x)$ is $x$, $A_{1}(x)$ is the parent of $x$, and so on. Fix a request $r$. For all sets of servers $S^{\prime},S^{\prime\prime}$ with $S^{\prime}\subseteq S^{\prime\prime}$, let $\ell^{\prime}$ satisfy $A_{\ell^{\prime}}(r)=A(r\mid S)$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime}$, and let $\ell^{\prime\prime}$ satisfy $A_{\ell^{\prime\prime}}(r)=A(r\mid S)$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}$. Note that $\ell^{\prime}$ and $\ell^{\prime\prime}$ only depend on $S^{\prime}$ and $S^{\prime\prime}$, respectively.
We first show that $\ell^{\prime}=\ell^{\prime\prime}$ by deriving a contradiction when assuming $\ell^{\prime}\neq\ell^{\prime\prime}$. We only analyze the case where $\ell^{\prime}<\ell^{\prime\prime}$, and the case where $\ell^{\prime}>\ell^{\prime\prime}$ can be proved analogously. When $\ell^{\prime}<\ell^{\prime\prime}$, all servers in $S^{\prime}$ are located in $T_{\ell^{\prime}}(r)$, and all servers in $S^{\prime\prime}$ are located in $T_{\ell^{\prime\prime}}(r)$. Moreover, by the definition of $\ell^{\prime\prime}$, no server in $S^{\prime\prime}$ is located in $T_{\ell^{\prime\prime}-1}(r)$. Since $T_{\ell^{\prime}}(r)$ is a subtree in $T_{\ell^{\prime\prime}-1}(r)$, we have $S^{\prime}\cap S^{\prime\prime}=\emptyset$, contradicting the assumption that $S^{\prime}\subseteq S^{\prime\prime}$.
Fix a server $s\in S^{\prime}$. For every $i\in[\ell^{\prime}]$, denote $w_{i}^{\prime}$ as the number of subtrees of $A_{i}(s)$ containing at least one available server upon the arrival of $r$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime}$. Moreover, define $w_{0}^{\prime}$ as the number of available servers collocated with $s$ (including $s$) upon the arrival of $r$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime}$. Recall that $\ell^{\prime}=\ell^{\prime\prime}$, and define $w_{i}^{\prime\prime}$ for all $i\in\{0,1,\ldots,\ell^{\prime}\}$ similarly when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}$. By the definition of the tie-breaking rule of ${\mathcal{A}}$,
@@
Since $S^{\prime}\subseteq S^{\prime\prime}$, we have $w_{i}^{\prime}\leq w_{i}^{\prime\prime}$ for all $i\in\{0,1,\ldots,\ell^{\prime}\}$. Therefore,
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime}\right\}\geq\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}\right\}\,,$
concluding that the tie-breaking rule of ${\mathcal{A}}$ is non-increasing.
Appendix C Stochastic Setting: Cost Functions Beyond Triangle Inequality
@@
In previous sections, we consider metric space where the cost function satisfies the exact triangle inequality. In this section, we showcase the generality of our reduction for the stochastic setting by proving positive results when the cost function only approximately satisfies the triangle inequality. Specifically, we say that a cost function $\delta$ on $X$ satisfies the $\eta$-approximate triangle inequality for $\eta>0$ if for all $x,y,z\in X$,
@@
We say that a metric space satisfies the $\eta$-approximate triangle inequality if the corresponding cost function satisfies the $\eta$-approximate triangle inequality.
We first generalize Theorem 2 to the case where the metric space only satisfies the approximate triangle inequality.
@@
Given a distribution $\mathbb{D}$ on a metric space $(X,\delta)$ satisfying the $\eta$-approximate triangle inequality for $\eta>0$, assume the existence of an algorithm ${\mathcal{A}}$ such that $\mathrm{cost}_{{\mathcal{A}}}(n)\leq\alpha\cdot\mathrm{OPT}(n)$. Then, an algorithm ${\mathcal{A}}^{\prime}$ exists such that $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\eta\cdot(\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n))$ for every set of servers $S$ with $|S|\geq n$. If we further have $|S|=n$, then $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq(2\alpha\eta^{2}+\eta)\cdot\mathrm{OPT}(S,n)$.
The proof of Theorem 10 can be obtained by replacing every usage of the triangle inequality by the approximate triangle inequality in the proofs of Theorem 1 and Theorem 2 and hence is omitted.
Next, we consider the case where $X=[0,1]^{d}$ and $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ for $p\geq 1$, [^6] Here, $p$ can be a real number. which we will prove that satisfies the $2^{p-1}$-approximate triangle inequality, and assume $\mathbb{D}$ to be uniform distribution. We show that in balanced markets with adversarial servers and stochastic requests, an $O(1)$-competitive algorithm exists for $d\geq 1$ and $p\in[1,\max\{d,2\})$.
Theorem 11.
Assume that $X=[0,1]^{d}$ and $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ for $d\geq 1$ and $p\in[1,\max\{d,2\})$. Supposing that $m=n$ and $\mathbb{D}$ is uniform distribution, an $O(1)$-competitive algorithm exists with adversarial servers and stochastic requests.
We remark here that competitive ratio guarantees for other combinations of $d$ and $p$, more general distributions, and unbalanced markets can be similarly achieved as in Section 4 and Section 5. Nevertheless, we didn’t pursue this direction since parts of the resulting competitive ratios seem suboptimal, i.e., super-constant, and no new algorithmic insight emerges.
C.1 Proof of Theorem 11
We start by showing that for all $p\geq 1$, the cost function $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ satisfies the $2^{p-1}$-approximate triangle inequality by proving a more general statement: Assuming $\delta^{\prime}$ to be a cost function satisfying the triangle inequality, the cost function defined as $\delta(x,y)\triangleq\delta^{\prime}(x,y)^{p}$ satisfies the $2^{p-1}$-approximate triangle inequality.
@@ C.1 Proof of Theorem 11
For every $p\geq 1$, given a cost function $\delta^{\prime}$ on $X$ satisfying the triangle inequality, the cost function on $X$ defined as $\delta(x,y)\triangleq\delta^{\prime}(x,y)^{p}$ for all $x,y\in X$ satisfies the $2^{p-1}$-approximate triangle inequality.
@@
where the first inequality holds by the triangle inequality and the second inequality holds by the convexity of $g(x)=x^{p}$ with $p\geq 1$. ∎
Next, we present the following estimation of $\mathrm{OPT}(n)$.
@@
Let $X=[0,1]^{d}$ and $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ for $d,p\geq 1$. Assuming ${\mathbb{D}}$ to be uniform distribution, it holds that
$\mathrm{OPT}(n)=\begin{cases}\Theta(n^{1-\frac{p}{2}})&d=1\\ \Theta(n^{1-\frac{p}{2}}(\log n)^{\frac{p}{2}})&d=2\\ \Theta(n^{1-\frac{p}{d}})&d\geq 3\end{cases}\,.$
Now, we are ready to prove Theorem 11. Let ${\mathcal{A}}$ be the algorithm as per Lemma 5, and we analyze $\mathrm{cost}_{{\mathcal{A}}}(n)$ for $d=1$, $d=2$, and $d\geq 3$ separately by applying Lemma 17. For $d=1$ and $p\in[1,2)$,
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}t^{-\frac{p}{2}}\right)=O\left(\int_{1}^{n}x^{-\frac{p}{2}}\mathrm{d}x\right)=O\left(n^{1-\frac{p}{2}}\right)=O\left(\mathrm{OPT}(n)\right)\,.$
@@
$\mathrm{cost}_{{\mathcal{A}}}(n)$ $=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}t^{-\frac{p}{2}}(\log t)^{\frac{p}{2}}\right)=O\left((\log n)^{\frac{p}{2}}\int_{1}^{n}x^{-\frac{p}{2}}\mathrm{d}x\right)$
$=O\left(n^{1-\frac{p}{2}}(\log n)^{\frac{p}{2}}\right)=O\left(\mathrm{OPT}(n)\right)\,.$
@@
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}t^{-\frac{p}{d}}\right)=O\left(\int_{1}^{n}x^{-\frac{p}{d}}\mathrm{d}x\right)=O\left(n^{1-\frac{p}{d}}\right)=O\left(\mathrm{OPT}(n)\right)\,.$
As a result, ${\mathcal{A}}$ admits an $O(1)$ competitive ratio when both servers and requests are uniformly distributed. By Lemma 16, $\delta$ satisfies the $2^{p-1}$-approximate triangle inequality. Combined with Theorem 10, we conclude that an $O(1)$-competitive algorithm exists in balanced markets with adversarial servers and uniformly distributed requests.
Appendix D Stochastic Setting: An Alternative Algorithm for Tree Metrics
In this section, we consider balanced markets and apply our approach to tree metrics with general distributions. In particular, a tree metric is defined by shortest path distances in a tree $T=(V,E)$. We will prove that the $\mathsf{SOAR}$ algorithm admits an $O(1)$ competitive ratio with all servers and requests independently drawn from $\mathbb{D}$, which implies the existence of an $O(1)$-competitive algorithm with servers adversarially chosen and requests independently drawn from $\mathbb{D}$ by Theorem 2. This gives an alternative $O(1)$-competitive algorithm in balanced markets with tree metrics and general distributions, recovering the result by [GGPW19].
We first present our main theorem in this section.
Theorem 12.
Suppose that $m=n$ and consider tree metrics. For an arbitrary distribution $\mathbb{D}$, an $O(1)$-competitive algorithm exists when servers are adversarially chosen and requests are independently drawn from $\mathbb{D}$.
To prove Theorem 12, it is known that in balanced markets with adversarial servers and stochastic requests following an arbitrary distribution, only a constant factor in the competitive ratio is lost by assuming that each request is uniformly distributed over server locations. That is, for every server $s\in S$, the request is collocated with $s$ with probability $1/n$.
Lemma 18 ([GGPW19]).
Suppose that $m=n$ and servers are adversarially chosen. For an arbitrary metric, if an $\alpha$-competitive algorithm exists when requests are uniformly distributed over server locations, then we can construct a $(2\alpha+1)$-competitive algorithm when requests follow an arbitrary distribution.
Given Lemma 18, it suffices to consider distribution $\mathbb{D}$ of the following form: Given a tree metric $T=(V,E)$, each node $v\in V$ is associated with an integer $a_{v}\geq 0$ such that $\sum_{v\in V}a_{v}=n$ and $\mathbb{P}_{r\sim\mathbb{D}}[r=v]=a_{v}/n$. For each edge $e\in E$, let $d_{e}$ denote the length of $e$, and let $T_{1}(e)$ and $T_{2}(e)$ be two subtrees of $T$ obtained by deleting $e$. For each $e\in E$, define $n_{e}=\sum_{v\in T_{1}(e)}a_{v}$. Given such a distribution $\mathbb{D}$, we estimate $\mathrm{OPT}(t)$ in the following lemma.
Lemma 19.
For each $t\in[n]$, $\mathrm{OPT}(t)=\Theta\left(\frac{\sqrt{t}}{n}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)$.
Proof.
Fix $t\in[n]$. For each $e\in E$, let $P_{e},Q_{e}\sim{\rm Binom}(t,n_{e}/n)$ be random variables denoting the number of servers located in $T_{1}(e)$ and the number of requests located in $T_{1}(e)$, respectively. Given a realization of $(P_{e},Q_{e})_{e\in E}$, in the perfect matching between servers and requests with the minimum cost, for each $e\in E$, exactly $|P_{e}-Q_{e}|$ matched server-request pairs will utilize $e$, incurring a cost of $d_{e}\cdot|P_{e}-Q_{e}|$. Hence,
$\mathrm{OPT}(t)=\mathbb{E}\left[\sum_{e\in E}d_{e}\cdot|P_{e}-Q_{e}|\right]=\sum_{e\in E}d_{e}\cdot\mathbb{E}[|P_{e}-Q_{e}|]\,.$
We first lower bound $\mathrm{OPT}(t)$, which relies on the following probabilistic bound.
Claim 1 ([BK13]).
Let $Z\sim{\rm Binom}(n,p)$, with $n\geq 2$ and $p\in[1/n,1-1/n]$. Then, we have
$\mathbb{E}[|Z-\mathbb{E}Z|]\geq\mathrm{std}(Z)/\sqrt{2}\,,$
where $\mathrm{std}(Z)$ denotes the standard deviation of $Z$.
We lower bound $\mathrm{OPT}(t)$ by
$\mathrm{OPT}(t)$ $=\sum_{e\in E}d_{e}\cdot\mathbb{E}[|P_{e}-Q_{e}|]\geq\sum_{e\in E}d_{e}\cdot\mathbb{E}[|P_{e}-\mathbb{E}P_{e}|]$
$\geq\frac{1}{\sqrt{2}}\sum_{e\in E}d_{e}\cdot\mathrm{std}(P_{e})=\Omega\left(\frac{\sqrt{t}}{n}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)\,,$
where the first inequality holds since the function $f(z)=\mathbb{E}[|P_{e}-z|]$ reaches its minimum when $z=\mathbb{E}P_{e}$, and the second inequality holds by Claim 1. Moreover, $\mathrm{OPT}(t)$ is upper bounded by
$\mathrm{OPT}(t)$ $=\sum_{e\in E}d_{e}\cdot\mathbb{E}[|P_{e}-Q_{e}|]\leq\sum_{e\in E}d_{e}\left(\mathbb{E}\left[\left|P_{e}-\mathbb{E}P_{e}\right|\right]+\mathbb{E}\left[\left|Q_{e}-\mathbb{E}Q_{e}\right|\right]\right)$
$\leq\sum_{e\in E}d_{e}(\mathrm{std}(P_{e})+\mathrm{std}(Q_{e}))=O\left(\frac{\sqrt{t}}{n}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)\,,$
where the first inequality holds by $\mathbb{E}P_{e}=\mathbb{E}Q_{e}$ for every $e\in E$. ∎
Now, we are ready to prove Theorem 12.
Proof of Theorem 12.
According to Lemma 18, we can assume requests to be uniformly distributed over server locations by losing a constant factor in the competitive ratio. In other words, let $a_{v}$ be the number of servers located at $v\in V$, the distribution $\mathbb{D}$ satisfies $\mathbb{P}_{r\sim\mathbb{D}}[r=v]=a_{v}/n$ for every $v\in V$. By Theorem 2, it suffices to give an $O(1)$-competitive algorithm when all servers and requests are independently drawn from $\mathbb{D}$.
Let ${\mathcal{A}}$ be the algorithm as per Lemma 5. Recall that $\mathrm{OPT}(t)$ is estimated by Lemma 19. When all servers and requests are independently drawn from $\mathbb{D}$, the expected cost of ${\mathcal{A}}$ equals
$\mathrm{cost}_{{\mathcal{A}}}(n)$ $=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}\frac{1}{n\sqrt{t}}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)$
$=O\left(\frac{1}{n}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\sum_{t=1}^{n}\frac{1}{\sqrt{t}}\right)=O\left(\frac{1}{\sqrt{n}}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)=O\left(\mathrm{OPT}(n)\right)\,,$
concluding that ${\mathcal{A}}$ is $O(1)$-competitive. ∎

### v3

@@ B.2 Proof of Lemma 13
We use $\mathcal{A}$} to denote the greedy subroutine in the Random Subtree algorithm. The tie-breaking rule of $\mathcal{A}$} is local by definition, and it remains to show that it is non-increasing. For all leaf node $x$ of the HST and $\ell\geq 0$, denote $A_{\ell}(x)$ as the $\ell$-th ancestor of $x$ and $T_{\ell}(x)$ as the subtree rooted at $A_{\ell}(x)$. That is, $A_{0}(x)$ is $x$, $A_{1}(x)$ is the parent of $x$, and so on. Fix a request $r$. For all sets of servers $S^{\prime},S^{\prime\prime}$ with $S^{\prime}\subseteq S^{\prime\prime}$, let $\ell^{\prime}$ satisfy $A_{\ell^{\prime}}(r)=A(r\mid S)$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime}$, and let $\ell^{\prime\prime}$ satisfy $A_{\ell^{\prime\prime}}(r)=A(r\mid S)$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}$. Note that $\ell^{\prime}$ and $\ell^{\prime\prime}$ only depend on $S^{\prime}$ and $S^{\prime\prime}$, respectively.
We first show that $\ell^{\prime}=\ell^{\prime\prime}$. Assume for contradiction that $\ell^{\prime}\neq\ell^{\prime\prime}$. We only analyze the case where $\ell^{\prime}<\ell^{\prime\prime}$, and the case where $\ell^{\prime}>\ell^{\prime\prime}$ can be handled analogously. When $\ell^{\prime}<\ell^{\prime\prime}$, all servers in $S^{\prime}$ are located in $T_{\ell^{\prime}}(r)$, and all servers in $S^{\prime\prime}$ are located in $T_{\ell^{\prime\prime}}(r)$. Moreover, by the definition of $\ell^{\prime\prime}$, no server in $S^{\prime\prime}$ is located in $T_{\ell^{\prime\prime}-1}(r)$. Since $T_{\ell^{\prime}}(r)$ is a subtree in $T_{\ell^{\prime\prime}-1}(r)$, we have $S^{\prime}\cap S^{\prime\prime}=\emptyset$, contradicting the assumption that $S^{\prime}\subseteq S^{\prime\prime}$.
Fix a server $s\in S^{\prime}$. For every $i\in[\ell^{\prime}]$, denote $w_{i}^{\prime}$ as the number of subtrees of $A_{i}(s)$ containing at least one available server upon the arrival of $r$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime}$. Moreover, define $w_{0}^{\prime}$ as the number of available servers collocated with $s$ (including $s$) upon the arrival of $r$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime}$. Recall that $\ell^{\prime}=\ell^{\prime\prime}$, and define $w_{i}^{\prime\prime}$ for every $i\in\{0,1,\ldots,\ell^{\prime}\}$ similarly when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}$. By the tie-breaking rule of $\mathcal{A}$},
@@
Since $S^{\prime}\subseteq S^{\prime\prime}$, we have $w_{i}^{\prime}\leq w_{i}^{\prime\prime}$ for every $i\in\{0,1,\ldots,\ell^{\prime}\}$. Therefore,
[…]
implying that the tie-breaking rule of $\mathcal{A}$} is non-increasing.
Appendix C Semi-Stochastic Setting: An Alternative $O(1)$-Competitive Algorithm for Tree Metrics with General Distributions
In this section, we consider balanced markets and tree metrics with general distributions. In particular, a tree metric is defined by the shortest path distances in a (weighted) tree $T=(V,E,w)$. We will prove that the SOAR algorithm given in Lemma 4 is $O(1)$-competitive in the stochastic setting, which leads to an $O(1)$-competitive algorithm in the semi-stochastic setting by Theorem 2, recovering the result by [GGPW19].
Theorem 10.
Suppose that $m=n$. Then, there exists an $O(1)$-competitive algorithm for the semi-stochastic setting and tree metrics with general distributions.
To prove Theorem 10, we utilize the result by [GGPW19], which holds for any metric, that only a constant factor in the competitive ratio is lost by assuming that $\mathbb{D}$} is the uniform distribution over server locations. That is, for every server $s\in S$, $\mathbb{P}_{r\sim{\mathbb{D}}}[r=s]=1/n$.
Lemma 15 (Lemma 2.1 in [GGPW19]).
Suppose that $m=n$ and consider the semi-stochastic setting. Given an $\alpha$-competitive algorithm for the uniform distribution over server locations, we can construct a $(2\alpha+1)$-competitive algorithm for general distributions.
Let $T=(V,E,w)$ be the tree that induces the tree metric. Given Lemma 15, it suffices to consider distribution $\mathbb{D}$ of the following form: Each node $v\in V$ is associated with an integer $a_{v}\geq 0$ such that $\sum_{v\in V}a_{v}=n$ and $\mathbb{P}_{r\sim\mathbb{D}}[r=v]=a_{v}/n$. For each edge $e\in E$, let $T_{1}(e)$ and $T_{2}(e)$ be two subtrees of $T$ obtained by deleting $e$, and define $n_{e}:=\sum_{v\in T_{1}(e)}a_{v}$. We give a tight characterization of $\mathrm{OPT}_{\mathbb{D}}(t)$ in the following lemma.
Lemma 16.
For each $t\in[n]$, $\mathrm{OPT}_{\mathbb{D}}(t)=\Theta\left(\frac{\sqrt{t}}{n}\sum_{e\in E}w_{e}\sqrt{n_{e}(n-n_{e})}\right)$.
Proof.
Fix $t\in[n]$. For each $e\in E$, let $P_{e}\sim{\rm Binom}(t,n_{e}/n)$ and $Q_{e}\sim{\rm Binom}(t,n_{e}/n)$ be random variables denoting the number of servers located in $T_{1}(e)$ and the number of requests located in $T_{1}(e)$, respectively. Given a realization of $(P_{e},Q_{e})_{e\in E}$, in the optimal perfect matching between servers and requests, for each edge $e\in E$, exactly $|P_{e}-Q_{e}|$ matched server-request pairs will utilize $e$ in their paths, incurring a cost of $w_{e}\cdot|P_{e}-Q_{e}|$. Hence,
$\mathrm{OPT}_{\mathbb{D}}(t)=\mathbb{E}\left[\sum_{e\in E}w_{e}\cdot|P_{e}-Q_{e}|\right]=\sum_{e\in E}w_{e}\cdot\mathbb{E}[|P_{e}-Q_{e}|]\,.$
The lower bound for $\mathrm{OPT}_{\mathbb{D}}(t)$ relies on the following probabilistic bound.
Lemma 17 ([BK13]).
Let $Z\sim{\rm Binom}(n,p)$, with $n\geq 2$ and $p\in[1/n,1-1/n]$. Then,
$\mathbb{E}[|Z-\mathbb{E}Z|]\geq\mathrm{std}(Z)/\sqrt{2}\,$,
where $\mathrm{std}(Z)$ denotes the standard deviation of $Z$.
We lower bound $\mathrm{OPT}_{\mathbb{D}}(t)$ by
$\mathrm{OPT}_{\mathbb{D}}(t)$ $=\sum_{e\in E}w_{e}\cdot\mathbb{E}[|P_{e}-Q_{e}|]\geq\sum_{e\in E}w_{e}\cdot\mathbb{E}[|P_{e}-\mathbb{E}P_{e}|]$
$\geq\frac{1}{\sqrt{2}}\sum_{e\in E}w_{e}\cdot\mathrm{std}(P_{e})=\Omega\left(\frac{\sqrt{t}}{n}\sum_{e\in E}w_{e}\sqrt{n_{e}(n-n_{e})}\right)\,$,
where the first inequality holds by Jensen’s inequality, and the second inequality holds by Lemma 17. Moreover, $\mathrm{OPT}_{\mathbb{D}}(t)$ is upper bounded by
$\mathrm{OPT}_{\mathbb{D}}(t)$ $=\sum_{e\in E}w_{e}\cdot\mathbb{E}[|P_{e}-Q_{e}|]\leq\sum_{e\in E}w_{e}\left(\mathbb{E}\left[\left|P_{e}-\mathbb{E}P_{e}\right|\right]+\mathbb{E}\left[\left|Q_{e}-\mathbb{E}Q_{e}\right|\right]\right)$
$\leq\sum_{e\in E}w_{e}(\mathrm{std}(P_{e})+\mathrm{std}(Q_{e}))=O\left(\frac{\sqrt{t}}{n}\sum_{e\in E}w_{e}\sqrt{n_{e}(n-n_{e})}\right)\,$,
where the first inequality holds since $\mathbb{E}P_{e}=\mathbb{E}Q_{e}$ for every $e\in E$, and the second inequality holds by Jensen’s inequality. ∎
Now, we are ready to prove Theorem 10.
Proof of Theorem 10.
We assume $\mathbb{D}$} to be the uniform distribution over server locations, which only causes an additional constant factor in the competitive ratio by Lemma 15. In other words, let $a_{v}$ be the number of servers located at $v\in V$, and it holds that $\mathbb{P}_{r\sim\mathbb{D}}[r=v]=a_{v}/n$. By Theorem 2, it suffices to give an $O(1)$-competitive algorithm in the stochastic setting.
Let $\mathcal{A}$} be the algorithm as per Lemma 4. By Lemma 16,
$\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)$ $=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}_{\mathbb{D}}(t)$
$=O\left(\sum_{t=1}^{n}\frac{1}{n\sqrt{t}}\sum_{e\in E}w_{e}\sqrt{n_{e}(n-n_{e})}\right)$
$=O\left(\frac{1}{n}\sum_{e\in E}w_{e}\sqrt{n_{e}(n-n_{e})}\sum_{t=1}^{n}\frac{1}{\sqrt{t}}\right)$
$=O\left(\frac{1}{\sqrt{n}}\sum_{e\in E}w_{e}\sqrt{n_{e}(n-n_{e})}\right)$
$=O\left(\mathrm{OPT}_{\mathbb{D}}(n)\right)\,$,
concluding that $\mathcal{A}$} is $O(1)$-competitive in the stochastic setting. ∎
Appendix D Semi-Stochastic Setting: Beyond Triangle Inequality
@@
In the remainder of the paper, we consider cost functions that satisfy the triangle inequality. In this section, we generalize some of our results to the case where the cost function only approximately satisfies the triangle inequality. Specifically, we say that a cost function $\delta:X\times X\to{\mathbb{R}}_{\geq 0}$ satisfies the $\eta$-approximate triangle inequality for $\eta\geq 1$ if for all $x,y,z\in X$,
@@
We first generalize Theorem 2 to the case where $\delta$ satisfies the approximate triangle inequality.
@@
Suppose that the cost function $\delta$ satisfies the $\eta$-approximate triangle inequality for $\eta\geq 1$. Then, for every distribution $\mathbb{D}$, given an algorithm $\mathcal{A}$} for online metric matching, there exists an algorithm $\mathcal{A}$$^{\prime}$ such that $\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}}^{\prime};S,n)\leq\eta\cdot(\mathrm{OPT}_{\mathbb{D}}(S,n)+\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n))$ for every set of servers $S$ with $|S|\geq n$. Moreover, for balanced markets, if $\mathcal{A}$} is $\alpha$-competitive in the stochastic setting, then $\mathcal{A}$$^{\prime}$ is $(2\alpha\eta^{2}+\eta)$-competitive in the semi-stochastic setting.
Theorem 11 is proved by replacing every usage of the triangle inequality by the approximate triangle inequality in the proofs of Theorems 1 and 2, and hence the proof is omitted.
We focus on the case where $X=[0,1]^{d}$ and $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ for $p\geq 1$, which we prove that satisfies the $2^{p-1}$-approximate triangle inequality. More generally, the following stronger statement holds: For $p\geq 1$, the cost function $\delta(x,y):=\delta^{\prime}(x,y)^{p}$, where $\delta^{\prime}$ satisfies the triangle inequality, satisfies the $2^{p-1}$-approximate triangle inequality.
@@
Suppose that $\delta:X\times X\to{\mathbb{R}}_{\geq 0}$ satisfies the triangle inequality. Then, the cost function $\delta(x,y):=\delta^{\prime}(x,y)^{p}$ for $p\geq 1$ satisfies the $2^{p-1}$-approximate triangle inequality.
@@
where the first inequality holds by the triangle inequality, and the second inequality holds by the convexity of $f(x)=x^{p}$ for $p\geq 1$. ∎
Assuming $\mathbb{D}$ to be the uniform distribution over $0,1]^{d}$, we give an $O(1)$-competitive algorithm for $d\geq 1$ and $p\in[1,\max\{d,2\})$ in the semi-stochastic setting.
Theorem 12.
Suppose that $m=n$ and $\mathbb{D}$ is the uniform distribution over $0,1]^{d}$. Let $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ for $p\in[1,\max\{d,2\})$. Then, for $d\geq 1$, there exists an $O(1)$-competitive algorithm in the semi-stochastic setting.
D.1 Proof of Theorem 12
We first recall the following estimate of $\mathrm{OPT}_{\mathbb{D}}(n)$.
@@ D.1 Proof of Theorem 12
Let $\mathbb{D}$} be the uniform distribution over $0,1]^{d}$, and let $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$. Then, for $d,p\geq 1$,
$\mathrm{OPT}_{\mathbb{D}}(n)=\begin{cases}\Theta(n^{1-\frac{p}{2}})&d=1\\ \Theta(n^{1-\frac{p}{2}}(\log n)^{\frac{p}{2}})&d=2\\ \Theta(n^{1-\frac{p}{d}})&d\geq 3\end{cases}\,.$
Now, we are ready to prove Theorem 12. Let $\mathcal{A}$} be the algorithm as per Lemma 4, and we analyze $\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)$ for $d=1$, $d=2$, and $d\geq 3$ separately by applying Lemma 19. For $d=1$ and $p\in[1,2)$,
$\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}_{\mathbb{D}}(t)=O\left(\sum_{t=1}^{n}t^{-\frac{p}{2}}\right)=O\left(\int_{1}^{n}x^{-\frac{p}{2}}\mathrm{d}x\right)=O\left(n^{1-\frac{p}{2}}\right)=O\left(\mathrm{OPT}_{\mathbb{D}}(n)\right)\,.$
@@
$\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)$ $=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}_{\mathbb{D}}(t)=O\left(\sum_{t=1}^{n}t^{-\frac{p}{2}}(\log t)^{\frac{p}{2}}\right)=O\left((\log n)^{\frac{p}{2}}\int_{1}^{n}x^{-\frac{p}{2}}\mathrm{d}x\right)$
$=O\left(n^{1-\frac{p}{2}}(\log n)^{\frac{p}{2}}\right)=O\left(\mathrm{OPT}_{\mathbb{D}}(n)\right)\,.$
@@
$\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}_{\mathbb{D}}(t)=O\left(\sum_{t=1}^{n}t^{-\frac{p}{d}}\right)=O\left(\int_{1}^{n}x^{-\frac{p}{d}}\mathrm{d}x\right)=O\left(n^{1-\frac{p}{d}}\right)=O\left(\mathrm{OPT}_{\mathbb{D}}(n)\right)\,.$
As a result, $\mathcal{A}$} is $O(1)$-competitive in the stochastic setting. By Lemma 18, $\delta$ satisfies the $2^{p-1}$-approximate triangle inequality. Combining with Theorem 11 leads to an $O(1)$-competitive algorithm in the semi-stochastic setting.